OpenStudy (anonymous):

Write an equation of the line through the given point with the given slope. Write the equation in slope-intercept form. (8, 0), m =1/2 A. y = 2x - 4 B. y= -1/2x-4 C. y= 1/2x-4 D. y= 1/2x+4

OpenStudy (anonymous):

y = mx + b y = 1/2x + b Plug in the coordinate: 0 = 1/2(8) + b 0 = 4 + b b = -4 Therefore: y = 1/2x - 4

OpenStudy (anonymous):

(y-0)=1/2(x-8) y=1/2x-4

OpenStudy (anonymous):

thanks everyone :)